# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 2.5 The branched, squashed three-sphere
2.6 The branched lens space
The three-dimensional orbifold encompassing the squashed three-sphere, the branched sphere and the squashed lens space is the branched lens space ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ with squashing parameters ${\left({b_{1},b_{2}}\right)}$ and conical singularities parametrized by the positive integers ${\left({{n_{+}},{n_{-}}}\right)}$. The branched lens space is a $\mathcal{O}(-n{n_{+}}{n_{-}})$ circle fibration over the spindle given by the weighted projective space $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$. Equivalently, ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ can be viewed as the free quotient $S^{3}_{b_{1},b_{2},{n_{+}},{n_{-}}}/\mathbb{Z}_{n}$ acting upon the Hopf fiber of the branched, squashed three-sphere. A line element on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ is
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\left({\frac{b_{1}}{{n_{+}}}}\right)}^{2}\sin^{2}x{\left({{\rm d}\varphi+\frac{{\rm d}\psi}{n}}\right)}^{2}+{\left({\frac{b_{2}}{n\,{n_{-}}}}\right)}^{2}\cos^{2}x\,{\rm d}\psi^{2}~{},$ (2.95)
where $\varphi,\psi\in{\left[{0,2\pi}\right]}$ are periodic and again $x\in{\left[{0,\pi/2}\right]}$. The function is the same $f$ used before, namely
$f=-\sqrt{b_{1}^{2}\cos^{2}x+b_{2}^{2}\sin^{2}x}~{}.$ (2.96)
From the line element above we see that the branched lens space reduces to the smooth (squashed) lens space if ${n_{+}}={n_{-}}=1$; while it becomes the branched-squashed three-sphere if $n=1$. Eventually, if ${n_{+}}={n_{-}}=n=1$, the branches lens space boils down to a squashed three-sphere.
In the orthonormal frame
$e^{1}$ $=-f\,{\rm d}x~{},\qquad e^{2}=b_{1}b_{2}\,\widetilde{f}^{-1}\sin x\cos x\,{\rm d}\varphi~{},\qquad e^{3}=\frac{\widetilde{f}}{n\,{n_{+}}{n_{-}}}{\left({{A^{(1)}}+{\rm d}\psi}\right)}~{},$
$\widetilde{f}$ $=\sqrt{{\left({{n_{-}}b_{1}}\right)}^{2}\sin^{2}x+{\left({{n_{+}}b_{2}}\right)}^{2}\cos^{2}x}~{},\qquad{A^{(1)}}=n{\left({\,{n_{-}}b_{1}\widetilde{f}^{-1}\sin x}\right)}^{2}{\rm d}\varphi~{},$ (2.97)
the branched lens space manifestly is a $\mathcal{O}{\left({-n\,{n_{+}}{n_{-}}}\right)}$ circle fibration over $\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ parametrized by ${\left({x,\varphi}\right)}$, indeed
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{{n_{+}}}+\frac{1}{{n_{-}}}=\chi_{\rm orb}~{},$ $\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=n=\frac{n\,{n_{+}}{n_{-}}}{{n_{+}}{n_{-}}}~{}.$ (2.98)
This geometry admits a Killing vector of the type (2.17), that is
$K={k_{0}}{\left({-\frac{{n_{-}}b_{1}+{n_{+}}b_{2}}{n\,{n_{-}}b_{1}}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ ${\omega}=-\frac{{n_{-}}b_{1}+{n_{+}}b_{2}}{n\,{n_{-}}b_{1}}~{},$ (2.99)
while its norm reads
${v}={k_{0}}b_{2}/{\left({n\,{n_{-}}}\right)}~{}.$ (2.100)
In turn, the Killing spinors (2.28), with components
$u_{1,2}$ $=\sqrt{\frac{{k_{0}}b_{2}}{2\,n\,{n_{-}}}{\left[{1\pm{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}\widetilde{f}^{-1}\sin x\cos x}\right]}}~{},$ (2.101)
@@ 2.6 The branched lens space
satisfy the Killing spinor equation for an arbitrary phase $\theta$ if the background fields
$A^{C}$ $={\rm d}\theta-\frac{b_{1}}{2{n_{+}}f}{\rm d}\varphi-\frac{{n_{-}}b_{1}-{n_{+}}b_{2}}{2n\,{n_{+}}{n_{-}}f}{\rm d}\psi~{},$
$V$ $=\frac{b_{1}{\left({1-{\rm i}fH}\right)}\sin^{2}x}{{n_{+}}f}{\rm d}\varphi+\frac{{\left[{{n_{-}}b_{1}-{n_{+}}b_{2}-{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}\cos{\left({2x}\right)}}\right]}{\left({1-{\rm i}fH}\right)}}{2n\,{n_{+}}{n_{-}}f}{\rm d}\psi~{},$
$A$ $=A^{C}+\frac{3}{2}V~{},$ (2.102)
are turned on. Such background fields are everywhere regular on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ if the phase of the Killing spinors $\theta$ is
$\alpha_{2}=-\frac{1}{{n_{+}}}~{},\qquad\alpha_{3}=\frac{{n_{+}}-{n_{-}}}{n\,{n_{+}}{n_{-}}}~{},\qquad\theta=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{}.$ (2.103)
@@ 3.1 BPS locus
${\mathcal{A}}^{N}={\mathcal{A}}^{N}_{{\left({0}\right)}}+\frac{\mathfrak{m}_{N}}{n_{N}}{\rm d}\varphi={\mathcal{A}}^{S}_{{\left({0}\right)}}+{\left({{\rm p}+\frac{\mathfrak{m}_{S}}{n_{S}}}\right)}{\rm d}\varphi={\mathcal{A}}^{S}+{\rm p}\,{\rm d}\varphi~{},$ (3.4)
@@
${\mathcal{A}}^{N}_{{\left({0}\right)}}$ $={\mathcal{A}}-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}{\rm d}\varphi~{},\qquad n_{N}={n_{+}}~{},\qquad\mathfrak{m}_{N}={\mathfrak{m}_{+}}~{},$
${\mathcal{A}}^{S}_{{\left({0}\right)}}$ $={\mathcal{A}}-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}{\rm d}\varphi~{},\qquad n_{S}={n_{-}}~{},\qquad\mathfrak{m}_{S}={\mathfrak{m}_{-}}-{n_{-}}{\rm p}~{}.$ (3.5)
@@
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{\mathcal{A}}={\rm p}+\frac{\mathfrak{m}_{S}}{n_{S}}-\frac{\mathfrak{m}_{N}}{n_{N}}=\frac{{\mathfrak{m}}}{{n_{+}}{n_{-}}}~{},$ (3.6)
@@
We do not impose any reality conditions on fields, for the moment. Inserting the gauge field (3.1) into the BPS equations above yields the BPS values of the scalar $\sigma$ and the auxiliary scalar $D$. Specifically, if ${\mathfrak{s}}=+1$, we have topologically twisted $\mathbb{\Sigma}\times S^{1}$ and the BPS locus
@@
whereas, if ${\mathfrak{s}}=-1$, we have anti-twisted $\mathbb{\Sigma}\times S^{1}$ and
@@ 3.2.1 Unpaired eigenvalues
$\mathfrak{p}_{+}=q_{G}{\mathfrak{m}_{+}}-{\mathfrak{s}}\frac{r}{2}~{},$ $\mathfrak{p}_{-}=q_{G}{\mathfrak{m}_{-}}+\frac{r}{2}~{},$
$\mathfrak{b}=1+{\mathfrak{s}}{\left\lfloor{{\mathfrak{s}}\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{},$ $\mathfrak{c}=\frac{{\left\llbracket{-\mathfrak{p}_{+}}\right\rrbracket_{{n_{+}}}}}{{n_{+}}}-{\mathfrak{s}}\frac{{\left\llbracket{{\mathfrak{s}}\,\mathfrak{p}_{-}}\right\rrbracket_{{n_{-}}}}}{{n_{-}}}~{},$
${\gamma_{R}}=-\frac{\alpha_{3}}{2}+\frac{{\omega}}{4}\chi_{-{\mathfrak{s}}}~{},$ ${\gamma_{G}}=-\varphi_{G}+\frac{{\omega}}{2}{\left({\frac{{\mathfrak{m}_{-}}}{{n_{-}}}+\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right)}~{},$
@@
where by ${\left\llbracket{\bullet}\right\rrbracket_{\diamond}}$ we indicate the reminder of the integer division of $\bullet$ by $\diamond$. In particular, ${\left({\mathfrak{b}-1}\right)}$ is the degree of the line orbibundle $\mathcal{O}(-\mathfrak{p})=\mathcal{O}(-{n_{+}}\mathfrak{p}_{-}+{n_{-}}\mathfrak{p}_{+})=\mathcal{O}{\left({-{\left({r/2}\right)}{\left({{n_{+}}+{n_{-}}}\right)}-q_{G}\mathfrak{m}}\right)}$, with $\mathfrak{p}={\left({{n_{+}}\mathfrak{p}_{-}-{n_{-}}\mathfrak{p}_{+}}\right)}$. Also, ${\omega}$, $\gamma_{G}$ and $\gamma_{R}$ are effective fugacities for angular momentum, gauge and R-symmetry, respectively. The objects written in (3.2.1) are valid for a theory invariant under a $U(1)$ gauge or flavour group. For a general, possibly non-Abelian, Lie group $G$, the substitutions
@@
should be applied to $Z^{\rm CM}_{\text{1-L}}$. The described substitutions correspond to the absence of regularity constraints on the eigenfunctions $\mathcal{B}_{m\sigma,m_{\psi}}$ and $\Phi_{n_{\sigma},n_{\psi}}$ at the north and south pole of the spindle. This approach becomes apparent when computing $Z^{\rm CM}_{\text{1-L}}$ through the index theorem, wherein calculations are performed separately in patches on $\mathbb{\Sigma}$. Notably, this method avoids generating absolute values in the one-loop determinants as demonstrated e.g. in [48].
@@
$Z^{\rm CM}_{\text{1-L}}=\prod_{\rho\in\mathfrak{R}_{G}}{\left({-y^{-\rho}}\right)}^{\frac{1}{4}{\left({2\mathfrak{b}-1+{\mathfrak{s}}}\right)}}q^{\frac{1}{8}{\left({1-{\mathfrak{s}}}\right)}{\left({\mathfrak{b}-1}\right)}}\frac{{\left({q^{\frac{1}{2}{\left({\mathfrak{b}+1}\right)}}y^{-\rho};q}\right)}_{\infty}}{{\left({q^{\frac{{\mathfrak{s}}}{2}{\left({\mathfrak{b}-1}\right)}}y^{-{\mathfrak{s}}\rho};q}\right)}_{\infty}}~{},$ (3.48)
@@
3.2.2 Equivariant orbifold index theorem

### v2

@@ 2.5 The branched-squashed three-sphere
2.6 The branched-squashed lens space
The squashed lens space and the branched-squashed three-sphere are specific instances of the branched-squashed lens space ${\mathbb{\Lambda}}$, which is described by a general $\mathcal{O}(-n)$ circle fibration over a spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$. A family of line elements on the three-dimensional orbifold ${\mathbb{\Lambda}}$ reads
${\rm d}s^{2}={\left({e^{1}}\right)}^{2}+{\left({e^{2}}\right)}^{2}+{\left({e^{3}}\right)}^{2}~{},$ (2.88)
with ${\left({e^{1},e^{2},e^{3}}\right)}$ being the orthonormal frame
$e^{1}$ $=-f\,{\rm d}x~{},$
$e^{2}$ $=b_{1}b_{2}\,\widetilde{f}^{-1}\sin x\cos x\,{\rm d}\varphi~{},$
$e^{3}$ $=\frac{\widetilde{f}}{n}{\left({{A^{(1)}}+{\rm d}\psi}\right)}~{},$
$\widetilde{f}$ $=\sqrt{{\left({{n_{-}}b_{1}\sin x}\right)}^{2}+{\left({{n_{+}}b_{2}\cos x}\right)}^{2}}~{},$
${A^{(1)}}$ $=n\,\widetilde{f}^{-2}{\left({{n_{-}}{t_{-}}b_{1}^{2}\sin^{2}x+{n_{+}}{t_{+}}b_{2}^{2}\cos^{2}x}\right)}{\rm d}\varphi~{},$ (2.89)
where the coordinates on the spindle base $\mathbb{\Sigma}$ are again ${\left({x,\varphi}\right)}$, the function $f$ is reported in (2.59) and ${\left({{t_{+}},{t_{-}}}\right)}$ are integers that satisfy the constraint ${\left({{n_{+}}{t_{-}}-{n_{-}}{t_{+}}}\right)}=1$ thanks to Bezout’s lemma. The Euler characteristic of the base and the flux of ${A^{(1)}}$ through $\mathbb{\Sigma}$ are
$\chi_{\rm orb}=\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{{n_{+}}}+\frac{1}{{n_{-}}}~{},$ $\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=\frac{n\,{t_{-}}}{{n_{-}}}-\frac{n\,{t_{+}}}{{n_{+}}}=\frac{n}{{n_{+}}{n_{-}}}~{},$ (2.90)
confirming that ${\mathbb{\Lambda}}$ is a $\mathcal{O}{\left({-n}\right)}$ circle bundle over $\mathbb{\Sigma}$. If $n=1$, then ${\mathbb{\Lambda}}$ coincides with a squashed three-sphere, which is indeed a $\mathcal{O}{\left({-1}\right)}$ circle fibration over a spindle. Furthermore, if $n={n_{+}}{n_{-}}$, then ${\mathbb{\Lambda}}$ becomes the branched-squashed three-sphere that we examined in a previous section. Moreover, if $n=n_{0}{n_{+}}{n_{-}}$ for some positive integer $n_{0}$, then ${\mathbb{\Lambda}}$ reduces to the free quotient $S^{3}_{b_{1},b_{2},{n_{+}},{n_{-}}}/\mathbb{Z}_{n_{0}}$, with $\mathbb{Z}_{n_{0}}$ acting upon the $S^{1}$ fiber of the branched-squashed three-sphere. For example, a line element on $S^{3}_{b_{1},b_{2},{n_{+}},{n_{-}}}/\mathbb{Z}_{n_{0}}$, is
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\left({\frac{b_{1}}{{n_{+}}}}\right)}^{2}\sin^{2}x{\left({{\rm d}\varphi+\frac{{\rm d}\psi}{n_{0}}}\right)}^{2}+{\left({\frac{b_{2}}{n_{0}\,{n_{-}}}}\right)}^{2}\cos^{2}x\,{\rm d}\psi^{2}~{},$ (2.91)
where, as usual, $\varphi,\psi\in{\left[{0,2\pi}\right]}$ are periodic and $x\in{\left[{0,\pi/2}\right]}$.
Let us now explore the geometric features of general ${\mathbb{\Lambda}}$ labelled by arbitrary integers ${\left({n,{n_{+}},{n_{-}},{t_{+}},{t_{-}}}\right)}$. Expanding (2.88) in a neighborhood of $x=0$ gives
${\rm d}s^{2}|_{x=0}=b_{1}^{2}{\left({{\rm d}x^{2}+x^{2}\frac{{\rm d}\varphi^{2}}{{n_{+}}^{2}}}\right)}+{\left({\frac{{n_{+}}b_{2}}{n}}\right)}^{2}{\left({{\rm d}\psi+\frac{n\,{t_{+}}}{{n_{+}}}{\rm d}\varphi}\right)}^{2}~{},$ (2.92)
which can be rewritten via the complex coordinates
$z_{0}^{(1)}=b_{1}\,x\,e^{{\rm i}\varphi/{n_{+}}}~{},$ $z_{0}^{(2)}=\frac{{n_{+}}b_{2}}{n}e^{{\rm i}{\left({\psi+\frac{n{t_{+}}}{{n_{+}}}\varphi}\right)}}~{},$ (2.93)
which are invariant under $\psi\sim\psi+2\pi$ and are subject to the identifications
$\varphi\sim\varphi+2\pi\implies{\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}\sim{\left({w_{+}z_{0}^{(1)},w_{+}^{n\,{t_{+}}}z_{0}^{(2)}}\right)}~{},$ $w_{+}=e^{2\pi{\rm i}/{n_{+}}}~{}.$ (2.94)
If ${\left\llbracket{n\,{t_{+}}}\right\rrbracket_{{n_{+}}}}=0$, where ${\left\llbracket{\bullet}\right\rrbracket_{\diamond}}$ is the reminder of the integer division of $\bullet$ by $\diamond$, then $w_{+}^{n\,{t_{+}}}=1$, implying that ${\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}$ are coordinates on the orbifold ${\left({\mathbb{C}/\mathbb{Z}_{{n_{+}}}}\right)}\times\mathbb{C}^{*}$, where $\mathbb{C}/\mathbb{Z}_{{n_{+}}}$ encodes a genuine conical singularity with deficit angle $2\pi{\left({1-{n_{+}}^{-1}}\right)}$. Otherwise, if $n\,{t_{+}}=k_{+}{t_{+}}^{\prime}$ and ${n_{+}}=k_{+}{n_{+}}^{\prime}$ share a common factor $k_{+}$, then winding ${n_{+}}^{\prime}$ times around the point $x=0$ shows that the north pole of the spindle base $\mathbb{\Sigma}$ is the fixed point of a $\mathbb{Z}_{k_{+}}$ action:
$\varphi\sim\varphi+2\pi\,{n_{+}}^{\prime}\implies{\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}\sim{\left({e^{2\pi{\rm i}/k_{+}}z_{0}^{(1)},z_{0}^{(2)}}\right)}~{}.$ (2.95)
Therefore, the topology of this fibration in a neighborhood of $x=0$ is that of ${\left({\mathbb{C}/\mathbb{Z}_{k_{+}}}\right)}\times\mathbb{C}^{*}$, with the quotient $\mathbb{C}/\mathbb{Z}_{k_{+}}$ describing the corresponding conical singularity. Eventually, if $n\,{t_{+}}$ and ${n_{+}}$ turn out to be mutually coprime, trivializing the $\mathbb{Z}_{{n_{+}}}$ action on the fiber also trivializes that on the base:
$\varphi\sim\varphi+2\pi{n_{+}}\implies{\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}\sim{\left({w_{+}^{{n_{+}}}z_{0}^{(1)},w_{+}^{n{n_{+}}{t_{+}}}z_{0}^{(2)}}\right)}={\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}~{}.$ (2.96)
In this case, the $\mathbb{Z}_{{n_{+}}}$ action is free as it exhibits no fixed points and the fibration is smooth near the point $x=0$. On the other hand,
${\rm d}s^{2}|_{x=\pi/2}=b_{2}^{2}{\left[{{\rm d}x^{2}+{\left({\frac{\pi}{2}-x}\right)}^{2}\frac{{\rm d}\varphi^{2}}{{n_{-}}^{2}}}\right]}+{\left({\frac{{n_{-}}b_{1}}{n}}\right)}^{2}{\left({{\rm d}\psi+\frac{n\,{t_{-}}}{{n_{-}}}{\rm d}\varphi}\right)}^{2}~{},$ (2.97)
suggesting the introduction of the complex coordinates
$z_{\pi/2}^{(1)}=b_{2}{\left({\frac{\pi}{2}-x}\right)}e^{-{\rm i}\varphi/{n_{-}}}~{},$ $z_{\pi/2}^{(2)}=\frac{{n_{-}}b_{1}}{n}e^{{\rm i}{\left({\psi+\frac{n{t_{-}}}{{n_{-}}}\varphi}\right)}}~{}.$ (2.98)
which are left unchanged by $\psi\sim\psi+2\pi$, while they fulfil the equivalence relations
$\varphi\sim\varphi+2\pi\implies{\left({z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}\sim{\left({w_{-}^{-1}z_{\pi/2}^{(1)},w_{-}^{-n\,{t_{-}}}z_{\pi/2}^{(2)}}\right)}~{},$ $w_{-}=e^{2\pi{\rm i}/{n_{-}}}~{},$ (2.99)
hold. Again, if ${\left\llbracket{n\,{t_{-}}}\right\rrbracket_{{n_{-}}}}=0$, the action on the fiber is trivial as $w_{-}^{n\,{t_{-}}}=1$ and the complex coordinates ${\left({z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}$ parametrize the orbifold ${\left({\mathbb{C}/\mathbb{Z}_{{n_{-}}}}\right)}\times\mathbb{C}^{*}$. Otherwise, if $n\,{t_{-}}=k_{-}{t_{-}}^{\prime}$ and ${n_{-}}=k_{-}{n_{-}}^{\prime}$, then the action on the fiber is trivialized by going ${n_{-}}^{\prime}$ times around the south pole of the spindle base, giving
$\varphi\sim\varphi+2\pi\,{n_{-}}^{\prime}\implies{\left({z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}\sim{\left({e^{-2\pi{\rm i}/k_{-}}z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}~{},$ (2.100)
implying that the topology of the fibration near the point $x=\pi/2$ is that of ${\left({\mathbb{C}/\mathbb{Z}_{k_{-}}}\right)}\times\mathbb{C}^{*}$. Finally, if $n\,{t_{-}}$ and ${n_{-}}$ are coprime, there is no way to trivialize the $\mathbb{Z}_{{n_{-}}}$ action on the fiber without trivializing its action on the base and $\mathbb{Z}_{{n_{-}}}$ has no fixed points. Hence, $\mathbb{Z}_{{n_{-}}}$ acts freely and the fibration is smooth in a neighborhood of $x=\pi/2$. If both $\mathbb{Z}_{{n_{+}}}$ and $\mathbb{Z}_{{n_{-}}}$ act freely everywhere, the fibration is a smooth manifold, as in the case of the lens space.
The Killing vector of the form (2.17) enjoyed by this fibration reads
$K=k_{0}{\left({{\omega}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ ${\omega}=-\frac{{n_{-}}b_{1}+{n_{+}}b_{2}}{n{\left({{t_{-}}b_{1}+{t_{+}}b_{2}}\right)}}~{},\qquad{v}=\frac{k_{0}b_{1}b_{2}}{n{\left({{t_{-}}b_{1}+{t_{+}}b_{2}}\right)}}~{},$ (2.101)
where ${v}={\left({\zeta\widetilde{\zeta}}\right)}$ is the contraction of the Killing spinors (2.1), with components
$u_{1,2}$ $=\sqrt{\frac{k_{0}\,b_{1}b_{2}}{2n{\left({{t_{-}}b_{1}+{t_{+}}b_{2}}\right)}}{\left[{1\pm{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}\widetilde{f}^{-1}\sin x\cos x}\right]}}~{},$ (2.102)
@@ 2.6 The branched-squashed lens space
and fulfilling the Killing spinor equation for an arbitrary phase $\theta$ if one turns on the background fields
$A^{C}$ $={\rm d}\theta+\frac{{t_{+}}b_{2}-{t_{-}}b_{1}}{2f}{\rm d}\varphi+\frac{{n_{+}}b_{2}-{n_{-}}b_{1}}{2nf}{\rm d}\psi~{},$
$V$ $={\left({{\rm i}H-\frac{1}{f}}\right)}{\left[{{\left({{t_{+}}b_{2}\cos^{2}x-{t_{-}}b_{1}\sin^{2}x}\right)}{\rm d}\varphi+{\left({{n_{+}}b_{2}\cos^{2}x-{n_{-}}b_{1}\sin^{2}x}\right)}\frac{{\rm d}\psi}{n}}\right]}~{},$
$A$ $=A^{C}+\frac{3}{2}V~{}.$ (2.103)
These background fields are regular on the fibration if
$\alpha_{2}={t_{+}}-{t_{-}}~{},\qquad\alpha_{3}=\frac{{n_{+}}-{n_{-}}}{n}~{},\qquad\theta=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{}.$ (2.104)
@@ 3.1 BPS locus
${\mathcal{A}}^{N}={\mathcal{A}}^{N}_{{\left({0}\right)}}+\frac{\mathfrak{m}_{N}}{{n_{+}}}{\rm d}\varphi={\mathcal{A}}^{S}_{{\left({0}\right)}}+{\left({{\rm p}+\frac{\mathfrak{m}_{S}}{{n_{-}}}}\right)}{\rm d}\varphi={\mathcal{A}}^{S}+{\rm p}\,{\rm d}\varphi~{},$ (3.4)
@@
${\mathcal{A}}^{N}_{{\left({0}\right)}}$ $={\mathcal{A}}-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}{\rm d}\varphi~{},\qquad{n_{+}}={n_{+}}~{},\qquad\mathfrak{m}_{N}={\mathfrak{m}_{+}}~{},$
${\mathcal{A}}^{S}_{{\left({0}\right)}}$ $={\mathcal{A}}-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}{\rm d}\varphi~{},\qquad{n_{-}}={n_{-}}~{},\qquad\mathfrak{m}_{S}={\mathfrak{m}_{-}}-{n_{-}}{\rm p}~{}.$ (3.5)
@@
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{\mathcal{A}}={\rm p}+\frac{\mathfrak{m}_{S}}{{n_{-}}}-\frac{\mathfrak{m}_{N}}{{n_{+}}}=\frac{{\mathfrak{m}}}{{n_{+}}{n_{-}}}~{},$ (3.6)
@@
We do not impose any reality conditions on fields, for the moment. Inserting the gauge field (3.1) into the BPS equations above yields the BPS values of the scalar $\sigma$ and the auxiliary scalar $D$. Specifically, if ${\upsigma}=+1$, we have topologically twisted $\mathbb{\Sigma}\times S^{1}$ and the BPS locus
@@
whereas, if ${\upsigma}=-1$, we have anti-twisted $\mathbb{\Sigma}\times S^{1}$ and
@@ 3.2.1 Unpaired eigenvalues
Here, we have set $\alpha_{2}=0$ for simplicity, in analogy with [36]. The identical final result follows for $\alpha_{2}\neq 0$ after consistently restoring $\alpha_{2}$ everywhere: for instance, if $\alpha_{2}\neq 0$, then (3.15) becomes
$\widehat{\Phi}_{n_{\varphi},n_{\psi}}=e^{{\rm i}{\left({r/2}\right)}\alpha_{2}+{\rm i}n_{\varphi}\varphi+{\rm i}n_{\psi}\psi}\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}~{},$ (3.16)
where $r$ is the $R$-charge of the chiral multiplet. The form of (3.16) ensures that $\widehat{\Phi}_{n_{\varphi},n_{\psi}}$ has the correct transition function on $\mathbb{\Sigma}$, as it behooves a section of the $R$-symmetry bundle on the spindle.
@@
$\mathfrak{p}_{+}=q_{G}{\mathfrak{m}_{+}}-{\upsigma}\frac{r}{2}~{},$ $\mathfrak{p}_{-}=q_{G}{\mathfrak{m}_{-}}+\frac{r}{2}~{},$
$\mathfrak{b}=1+{\upsigma}{\left\lfloor{{\upsigma}\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{},$ $\mathfrak{c}=\frac{{\left\llbracket{-\mathfrak{p}_{-}}\right\rrbracket_{{n_{-}}}}}{{n_{-}}}-{\upsigma}\frac{{\left\llbracket{{\upsigma}\,\mathfrak{p}_{+}}\right\rrbracket_{{n_{+}}}}}{{n_{+}}}~{},$
${\gamma_{R}}=-\frac{\alpha_{3}}{2}+\frac{{\omega}}{4}\chi_{-{\upsigma}}~{},$ ${\gamma_{G}}=-\varphi_{G}+\frac{{\omega}}{2}{\left({\frac{{\mathfrak{m}_{-}}}{{n_{-}}}+\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right)}~{},$
@@
In particular, ${\left({\mathfrak{b}-1}\right)}$ is the degree of the line bundle $\mathcal{O}(-\mathfrak{p})=\mathcal{O}{\left({-{\left({r/2}\right)}{\left({{n_{+}}+{n_{-}}}\right)}-q_{G}\mathfrak{m}}\right)}$, with $\mathfrak{p}={\left({{n_{+}}\mathfrak{p}_{-}-{n_{-}}\mathfrak{p}_{+}}\right)}$. Also, ${\omega}$, $\gamma_{G}$ and $\gamma_{R}$ are effective fugacities for angular momentum, gauge and $R$-symmetry, respectively. The objects written in (3.2.1) are valid for a theory invariant under a $U(1)$ gauge or flavour group. For a general, possibly non-Abelian, Lie group $G$, the substitutions
@@
should be applied to $Z^{\rm CM}_{\text{1-L}}$. The described substitutions correspond to the absence of regularity constraints on the eigenfunctions $\mathcal{B}_{m\sigma,m_{\psi}}$ and $\Phi_{n_{\sigma},n_{\psi}}$ at the north and south pole of the spindle. This approach becomes apparent when computing $Z^{\rm CM}_{\text{1-L}}$ through the index theorem, wherein calculations are performed separately in patches on $\mathbb{\Sigma}$. Notably, this method avoids generating absolute values in the one-loop determinants, as shown e.g. in [49].
@@
$Z^{\rm CM}_{\text{1-L}}=\prod_{\rho\in\mathfrak{R}_{G}}{\left({-y^{-\rho}}\right)}^{\frac{1}{4}{\left({2\mathfrak{b}-1+{\upsigma}}\right)}}q^{\frac{1}{8}{\left({1-{\upsigma}}\right)}{\left({\mathfrak{b}-1}\right)}}\frac{{\left({q^{\frac{1}{2}{\left({\mathfrak{b}+1}\right)}}y^{-\rho};q}\right)}_{\infty}}{{\left({q^{\frac{{\upsigma}}{2}{\left({\mathfrak{b}-1}\right)}}y^{-{\upsigma}\rho};q}\right)}_{\infty}}~{},$ (3.49)
@@
3.2.2 Equivariant index theorem on orbifolds